# Robust Bayesian Inference for Berkson and Classical Measurement Error Models

### v2

@@ B.1 Proofs of theoretical results for the Berkson error model
$\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0})$ $\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{XY})$
$\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{XY})$
$\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})$
$\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+2\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})$
@@
$\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta}^{0},\operatorname{\mathbb{P}}_{\theta})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})$
@@
where we used the triangle inequality in the first two and last two inequalities and the definition of $\theta^{*}(\operatorname{\mathbb{P}})$ in the third inequality. Using lemma B.1 with $S_{i}=(x_{i},y_{i})\sim Q_{i}:=\operatorname{\mathbb{P}}^{0}_{X\>|\>w_{i}}\operatorname{\mathbb{P}}_{Y\>|\>x}^{0}$ we have that:
@@
Using equation (38) and lemmas B.4 and B.5, we have that:
@@
$\quad\quad\leq\frac{4}{\sqrt{n}}+2(1+\Lambda)\left(\sqrt{\frac{1}{c+2}}+\frac{c}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N},\operatorname{\mathbb{F}}_{N}^{0})+\frac{1}{c+1}\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right).$
B.1.2 Proof of Corollary 4.2.1 (Berkson)
@@
B.1.3 Proof of Corollary 4.2.3 (Berkson)
@@
$\operatorname{\operatorname{MMD}}^{2}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N},\operatorname{\mathbb{F}}_{N}^{0})$ $=\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}k^{2}_{X}(x,x^{\prime})d\operatorname{\mathbb{F}}_{N}^{0}d\operatorname{\mathbb{F}}_{N}^{0}-2\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}k^{2}_{X}(x,x^{\prime})d\operatorname{\mathbb{F}}_{N}^{0}d\operatorname{\mathbb{F}}_{N}$
@@
$\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}k^{2}_{X}(x,x^{\prime})d\operatorname{\mathbb{F}}_{N}^{0}d\operatorname{\mathbb{F}}_{N}$ $=\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\exp\left\{-\frac{1}{l^{2}}\|x-x^{\prime}\|^{2}\right\}(2\pi\sigma_{1}^{2})^{-d_{\mathcal{X}}/2}\exp\left\{-\frac{1}{2\sigma_{1}^{2}}\|x\|^{2}\right\}$
@@
$=\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\exp\left\{-\frac{1}{2l^{\prime 2}}\|x-x^{\prime}\|^{2}\right\}(2\pi\sigma_{1}^{2})^{-d_{\mathcal{X}}/2}\exp\left\{-\frac{1}{2\sigma_{1}^{2}}\|x\|^{2}\right\}$
@@
$=(2\pi l^{\prime 2})^{d_{\mathcal{X}}/2}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\left(\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}(2\pi l^{\prime 2})^{-d_{\mathcal{X}}/2}\exp\left\{-\frac{1}{2l^{\prime 2}}\|x-x^{\prime}\|^{2}\right\}\right.$
@@
where in the second equality we rewrote $\exp\left\{-\frac{1}{2\sigma_{2}^{2}}\|x^{\prime}\|^{2}\right\}=\exp\left\{-\frac{1}{2\sigma_{2}^{2}}\|x^{\prime}-w\|^{2}\right\}$ which holds true in the case $w=0$. The inner integral is the density function of the sum of two Gaussian RVs with mean 00 and variances $l^{2}$ and $\sigma_{1}^{2}$ respectively. We will denote the sum of these RVs $Z=X+Y$ where $X\sim N(0,l^{2}I)$ and $Y\sim N(0,\sigma_{1}^{2}I)$ such that the inner integral corresponds to the probability density function of $Z$. Hence the above becomes
[…]
@@
[…]
$=(2\pi l^{\prime 2})^{d_{\mathcal{X}}/2}(2\pi(l^{\prime 2}+\sigma_{1}^{2}+\sigma_{2}^{2}))^{-d_{\mathcal{X}}/2}\exp\left\{-\frac{1}{2(l^{\prime 2}+\sigma_{1}^{2}+\sigma_{2}^{2})}\|0\|^{2}\right\}$
$=\left(\frac{l^{\prime 2}}{l^{\prime 2}+\sigma_{1}^{2}+\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}$
$=\left(\frac{l^{2}}{l^{2}+2\sigma_{1}^{2}+2\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}.$
We can use the same argument to derive the other two terms in (B.1.3) to obtain
@@
By Jensen’s inequality,
$\mathbb{E}_{\nu^{0}_{1:n}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}_{N}^{0}}\left[\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right]$ $\leq\sqrt{\mathbb{E}_{\nu^{0}_{1:n}\operatorname{\overset{iid}{\sim}}\operatorname{\mathbb{F}}_{N}^{0}}\left[\operatorname{\operatorname{MMD}}^{2}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})\right]}$
$=\left(\left(\frac{l^{2}}{l^{2}+4\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+2\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}+1\right)^{\frac{1}{2}}.$
@@
$\quad\quad\leq\frac{4}{\sqrt{n}}+2(1+\Lambda)\left(\sqrt{\frac{1}{c+2}}+C_{1}\frac{c}{c+1}+\frac{C_{2}}{c+1}\right)$
where $C_{1}^{2}:=\left(\frac{l^{2}}{l^{2}+4\sigma_{1}^{2}}\right)^{d_{\mathcal{X}}/2}-2\left(\frac{l^{2}}{l^{2}+2\sigma_{1}^{2}+2\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}+\left(\frac{l^{2}}{l^{2}+4\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}$ and $C_{2}^{2}:=\left(\frac{l^{2}}{l^{2}+4\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+2\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}+1$. ∎
B.2 Proofs of theoretical results for the Classical error model
B.2.1 Proof of Theorem 4.2 (Classical)

### v3

@@ D.1 Proofs of theoretical results for the Berkson error model
$\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0})$
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{XY})$
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})$
$\quad\quad+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{XY})$
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}_{XY})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})$
$\quad\quad+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0,n},\operatorname{\mathbb{P}}_{XY})$
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{XY},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})$
$\quad\quad+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0,n},\operatorname{\mathbb{P}}_{XY})$
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+2\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})$
@@
$\quad\leq\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0}_{\theta},\operatorname{\mathbb{P}}^{0})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta}^{0},\operatorname{\mathbb{P}}_{\theta})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})$
@@
where we used the triangle inequality in the first two and last three inequalities and the definition of $\theta^{\star}(\operatorname{\mathbb{P}})$ in the third inequality. Since this holds for any $\theta\in\Theta$ we have that
$\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0})$ $\leq\inf_{\theta\in\Theta}\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0}_{\theta})+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta}^{0},\operatorname{\mathbb{P}}_{\theta})$
$\quad+\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})}^{0},\operatorname{\mathbb{P}}_{\theta^{\star}(\operatorname{\mathbb{P}})})+2\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0},\operatorname{\mathbb{P}}^{0,n})$
$\quad+2\operatorname{\operatorname{MMD}}_{k}(\operatorname{\mathbb{P}}^{0,n},\operatorname{\mathbb{P}}_{XY}).$
Using lemma D.1 with $S_{i}=(x_{i},y_{i})\sim Q_{i}:=\operatorname{\mathbb{P}}^{0}_{X\>|\>w_{i}}\operatorname{\mathbb{P}}_{Y\>|\>x}^{0}$ we have that:
@@
Using (35) and Lemmas D.4 and D.5, we have that:
@@
$\leq\frac{1}{\sqrt{n}}\left(2+\frac{2(1+\Lambda)}{\sqrt{c+2}}\right)$
$\quad+\frac{2c}{c+1}\left(\Lambda\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N},\operatorname{\mathbb{F}}_{N}^{0})+\sqrt{{\text{Var}}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]+\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\operatorname{\mathbb{F}}_{N})}\right)$
$\quad+\frac{2}{c+1}\left(\Lambda\operatorname{\operatorname{MMD}}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})+\sqrt{2}\sqrt{\operatorname{\operatorname{MMD}}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})}\right).$
D.1.2 Proof of Corollary 4.2.1 (Berkson)
@@
D.1.3 Proof of Corollary 4.2.3 (Berkson)
@@
$\operatorname{\operatorname{MMD}}^{2}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N},\operatorname{\mathbb{F}}_{N}^{0})$
$\quad=\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}k^{2}_{X}(x,x^{\prime})d\operatorname{\mathbb{F}}_{N}^{0}d\operatorname{\mathbb{F}}_{N}^{0}-2\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}k^{2}_{X}(x,x^{\prime})d\operatorname{\mathbb{F}}_{N}^{0}d\operatorname{\mathbb{F}}_{N}$
@@
$\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}k^{2}_{X}(x,x^{\prime})d\operatorname{\mathbb{F}}_{N}^{0}d\operatorname{\mathbb{F}}_{N}$
$\quad=\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\exp\left\{-\frac{1}{l^{2}}\|x-x^{\prime}\|^{2}\right\}(2\pi\sigma_{1}^{2})^{-d_{\mathcal{X}}/2}\exp\left\{-\frac{1}{2\sigma_{1}^{2}}\|x\|^{2}\right\}$
@@
$\quad=\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\exp\left\{-\frac{1}{2l^{\prime 2}}\|x-x^{\prime}\|^{2}\right\}(2\pi\sigma_{1}^{2})^{-d_{\mathcal{X}}/2}\exp\left\{-\frac{1}{2\sigma_{1}^{2}}\|x\|^{2}\right\}$
@@
$\quad=(2\pi l^{\prime 2})^{d_{\mathcal{X}}/2}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\left(\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}(2\pi l^{\prime 2})^{-d_{\mathcal{X}}/2}\exp\left\{-\frac{1}{2l^{\prime 2}}\|x-x^{\prime}\|^{2}\right\}\right$.
@@
where in the second equality we rewrote
$\exp\left\{-\frac{1}{2\sigma_{2}^{2}}\|x^{\prime}\|^{2}\right\}=\exp\left\{-\frac{1}{2\sigma_{2}^{2}}\|x^{\prime}-w\|^{2}\right\}$
which holds true in the case $w=0$. The inner integral is the density function of the sum of two Gaussian RVs with mean 0 and variances $l^{2}$ and $\sigma_{1}^{2}$ respectively. We will denote the sum of these RVs $Z=X+Y$ where $X\sim N(0,l^{2}I)$ and $Y\sim N(0,\sigma_{1}^{2}I)$ such that the inner integral corresponds to the probability density function of $Z$. Hence the above becomes
$\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}k^{2}_{X}(x,x^{\prime})d\operatorname{\mathbb{F}}_{N}^{0}d\operatorname{\mathbb{F}}_{N}$
$\quad=(2\pi l^{\prime 2})^{d_{\mathcal{X}}/2}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}f_{Z}(x^{\prime})(2\pi\sigma_{2}^{2})^{-d_{\mathcal{X}}/2}\exp\left\{-\frac{1}{2\sigma_{2}^{2}}\|x^{\prime}-w\|^{2}\right\}dx^{\prime}$
@@
$\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}k^{2}_{X}(x,x^{\prime})d\operatorname{\mathbb{F}}_{N}^{0}d\operatorname{\mathbb{F}}_{N}$
$\quad=(2\pi l^{\prime 2})^{d_{\mathcal{X}}/2}\int_{\operatorname{\mathbb{R}}^{d_{\mathcal{X}}}}f_{Z}(x^{\prime})(2\pi\sigma_{2}^{2})^{-d_{\mathcal{X}}/2}exp\left\{-\frac{1}{2\sigma_{2}^{2}}\|x^{\prime}-w\|^{2}\right\}dx^{\prime}$
$\quad=(2\pi l^{\prime 2})^{d_{\mathcal{X}}/2}(2\pi(l^{\prime 2}+\sigma_{1}^{2}+\sigma_{2}^{2}))^{-d_{\mathcal{X}}/2}\exp\left\{-\frac{1}{2(l^{\prime 2}+\sigma_{1}^{2}+\sigma_{2}^{2})}\|0\|^{2}\right\}$
$\quad=\left(\frac{l^{\prime 2}}{l^{\prime 2}+\sigma_{1}^{2}+\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}$
$\quad=\left(\frac{l^{2}}{l^{2}+2\sigma_{1}^{2}+2\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}.$
We can use the same argument to derive the other two terms in (D.1.3) to obtain
@@
Note that using exactly the same argument, we obtain:
$\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{F}}_{N},\operatorname{\mathbb{F}}_{N}^{0})=\left(\frac{l^{2}}{l^{2}+2\sigma_{1}^{2}}\right)^{d_{\mathcal{X}}/2}-2\left(\frac{l^{2}}{l^{2}+\sigma_{1}^{2}+\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}+\left(\frac{l^{2}}{l^{2}+2\sigma_{2}^{2}}\right)^{d_{\mathcal{X}}/2}.$
@@
and
$\operatorname{\operatorname{MMD}}^{2}_{k_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})=\left(\frac{l^{2}}{l^{2}+2\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}+1.$
For the variance term, we have:
$\text{Var}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]$ $=\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[k_{X}(\nu,\nu)]-\left<\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[\phi_{X}(\nu)],\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0}}[\phi_{X}(\nu)]\right$>
$=1-\operatorname{\mathbb{E}}_{\nu\sim\operatorname{\mathbb{F}}_{N}^{0},\nu^{\prime}\sim\operatorname{\mathbb{F}}_{N}^{0}}[k_{X}(\nu,\nu^{\prime})]$
$=1-\left(\frac{l^{2}}{l^{2}+2\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}$
and hence
$\operatorname{\operatorname{MMD}}^{2}_{k^{2}_{X}}(\operatorname{\mathbb{F}}_{N}^{0},\delta_{0})+\text{Var}_{\mathcal{H}_{k_{X}}}[\operatorname{\mathbb{F}}_{N}^{0}]$
$\quad=1-2\left(\frac{l^{2}}{l^{2}+\sigma_{1}^{2}+\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}+\left(\frac{l^{2}}{l^{2}+2\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}.$
@@
$\quad\quad\leq\frac{1}{\sqrt{n}}\left(2+\frac{2(1+\Lambda)}{\sqrt{c+2}}\right)+\frac{2c}{c+1}\left(\Lambda C_{1}+C_{3}\right)+\frac{2}{c+1}\left(\Lambda C_{2}+\sqrt{2}\sqrt{C_{4}}\right).$
where
$C_{1}^{2}$ $:=\left(\frac{l^{2}}{l^{2}+4\sigma_{1}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+2\sigma_{1}^{2}+2\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}+\left(\frac{l^{2}}{l^{2}+4\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}$
$C_{2}^{2}$ $:=\left(\frac{l^{2}}{l^{2}+4\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+2\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}+1$
$C_{3}^{2}$ $:=1-2\left(\frac{l^{2}}{l^{2}+\sigma_{1}^{2}+\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}+\left(\frac{l^{2}}{l^{2}+2\sigma_{2}^{2}}\right)^{\frac{d_{\mathcal{X}}}{2}}$
$C_{4}^{2}$ $:=\left(\frac{l^{2}}{l^{2}+2\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}-2\left(\frac{l^{2}}{l^{2}+\sigma^{2}_{1}}\right)^{\frac{d_{\mathcal{X}}}{2}}+1.$
D.2 Proofs of theoretical results for the Classical error model
D.2.1 Proof of Theorem 4.2 (Classical)